OpenStudy (collegekid):

solve for x and show check. -4(x-2)+8=9x+3

OpenStudy (anonymous):

so... -4(x-2)+8=9x+3 multiplying the parentheses turns into -4x+8+8=9x+3 Combine common variables. you get -4x+16=9x+3. add the 4x over. 16=13x+3 subtract the 3 13=13x. divide by 13. x=1